### v1

# Searching for leptophilic composite asymmetric dark sector at $e^{+}e^{-}$ colliders

Abstract
Composite asymmetric dark matter (ADM) models provide a well-motivated paradigm that simultaneously explains dark matter (DM) relic density and matter-antimatter asymmetry. In these models, the mass of the DM candidate (the lightest dark baryon) is generated through the dark confinement scale dynamics. Although the leptophilic composite ADM model offers a viable framework, comprehensive studies of its collider phenomenology are absent. This work systematically explores novel signatures from leptophilic composite asymmetric dark sector at both low-energy and high-energy $e^{+}e^{-}$ colliders as well as other existing collider constraints. We demonstrate detectability of TeV-scale mediators along with sub-GeV to GeV-scale lightest dark mesons at Belle II and its proposed far detector, GAZELLE, as well as CEPC experiments. Moreover, these experiments exhibit complementary coverage of the model parameter space.
preprint: CPTNP-2025-034
I Introduction
The particle nature of dark matter (DM) and the origin of matter-antimatter asymmetry remain outstanding challenges in particle physics, astrophysics, and cosmology. The proposed DM models span mass scales from ultralight to tens of solar masses Battaglieri:2017aum ; Lin:2019uvt ; Ferreira:2020fam ; Belenchia:2021rfb , with well-motivated candidates including: weakly interacting massive particles (WIMPs) Arcadi:2017kky ; Arcadi:2024ukq , strongly interacting massive particles (SIMPs) Hochberg:2014dra , hidden sector DM Battaglieri:2017aum , WIMPless DM Feng:2008ya , asymmetric DM (ADM) Petraki:2013wwa ; Zurek:2013wia , freeze-in DM Bernal:2017kxu , QCD axions Baryakhtar:2025jwh , and primordial black holes Villanueva-Domingo:2021spv . Despite extensive searches, no conclusive evidence supports any candidate to date. Consequently, diverse experimental strategies have been developed to probe their properties.
Dynamically generating the observed matter-antimatter asymmetry requires satisfying Sakharov’s three conditions Sakharov:1967dj : 1) baryon number violation, 2) $C/CP$ violation, and 3) departure from thermal equilibrium. Successful mechanisms include electroweak baryogenesis Bodeker:2020ghk , leptogenesis Davidson:2008bu , and the Affleck-Dine mechanism Affleck:1984fy . The observed DM-to-baryon density ratio $\Omega_{\text{DM}}/\Omega_{\text{B}}\approx 5$ suggests a common origin for both abundances. The ADM paradigm Petraki:2013wwa ; Zurek:2013wia achieves this by positing that DM abundance, like baryonic abundance, arises from particle-antiparticle asymmetry generated by shared physics connecting dark and Standard Model (SM) sectors.
Composite ADM Blinnikov:1982eh ; Blinnikov:1983gh ; Khlopov:1989fj ; Alves:2009nf ; SpierMoreiraAlves:2010err ; Bai:2013xga ; Ibe:2018juk ; Ibe:2018tex ; Ibe:2019ena ; Hall:2019rld ; Zhang:2021orr ; Bottaro:2021aal ; Ibe:2021gil ; Ritter:2022opo represents a prominent scenario in which the dark sector features a hidden $SU(N)$ gauge symmetry analogous to SM QCD. In such models, accidental baryon number conservation stabilizes the lightest dark baryon, making it a natural DM candidate. Crucially, the DM mass is directly related to the dark confinement scale $\Lambda_{d}$, thereby being dynamically determined rather than a free parameter. Mediators facilitate dark sector and SM sector coupling to generate shared particle-antiparticle asymmetry. We adopt the framework in Refs. Bai:2013xga ; Zhang:2021orr where bi-charged scalar mediator decays produce asymmetries in both sectors.
Two bi-charged scalar mediators exist in the literature: baryophilic Bai:2013xga and leptophilic Zhang:2021orr types, both capable of simultaneously explaining DM relic density and matter-antimatter asymmetry. The baryophilic mediator carries charges under both hidden $SU(N)$ and SM QCD symmetries, coupling to a dark quark and an SM quark. This enables copious dark quark production at the LHC where, given the dark quark mass is much smaller than the collision energy, the produced dark quarks become energetic and highly boosted. When the scale of $\Lambda_{d}$ is also much smaller than the collision energy, subsequent QCD-like showering and hadronization generate distinctive dark jets that facilitate novel search strategies at the large hadron collider (LHC) Schwaller:2015gea ; Cohen:2015toa ; Cohen:2017pzm ; Park:2017rfb ; Renner:2018fhh ; Mies:2020mzw ; Linthorne:2021oiz ; Archer-Smith:2021ntx ; Carrasco:2023loy ; Carmona:2024tkg . In contrast, the leptophilic mediator is coupled with a dark quark and an SM lepton, suppressing LHC production rates as noted in Zhang:2021orr . Therefore, lepton colliders consequently provide ideal environments for probing such models. Although Ref. Zhang:2021orr proposed preliminary studies, comprehensive analyzes remain lacking. This work therefore systematically explores leptophilic composite ADM signatures on low-energy and high-energy $e^{+}e^{-}$ colliders.
Collider searches for leptophilic scalar mediators resemble searches of slepton in supersymmetry models when dark hadrons inside dark jets escape detection at LHC and LEP ATLAS:2019lff ; CMS:2020bfa ; CMS:2022syk ; CMS:2024gyw ; ATLAS:2024fub ; Fox:2011fx . We therefore rescale existing slepton constraints to these mediators. While collision energies exceeding the dark confinement scale $\Lambda_{d}$ can produce dark showers, the intensity of showering depends critically on energy: higher energies generate denser parton showers. At low-energy $e^{+}e^{-}$ colliders such as $B$-factories, sparse dark showers enable tracking of individual displaced dark pions and kinematic reconstruction. Similar approaches have been studied for the dark photon mediator Bernreuther:2022jlj , though signatures differ fundamentally from this work. Conversely, at high-energy $e^{+}e^{-}$ colliders (CEPC, FCC-ee, ILC), dense dark showers necessitate identification of displaced, collimated lepton jets rather than individual dark pions Zhang:2021orr . When dark jets become undetectable, mono-photon signatures Liu:2019ogn become relevant. This work establishes complementary coverage of parameter space through these distinct search strategies.
This paper is structured as follows: Section II reviews simplified leptophilic composite ADM models. Section III examines existing collider constraints. Novel dark shower signatures at $B$-factories and CEPC are explored in Sections IV and V, respectively. Section VI presents our conclusions and future research directions.
II A brief review of QCD-like dark sector with bi-charged mediators
In the leptophilic composite ADM models proposed in Ref. Zhang:2021orr , new heavy particles, including Majorana fermions and $SU(3)_{D}$ charged Dirac fermions, generate out-of-equilibrium processes and explicitly break hidden baryon number conservation. However, these particles lie significantly above current collider energy scales and are consequently decoupled from our analysis. We therefore restrict the Lagrangian to the relevant scalar mediator $X$ and dark quark $q_{d}$ interactions as follows [^1] The possible Higgs portal coupling between the SM Higgs doublet field and the $X$ field is irrelevant to this study. Therefore, we assume that its contributions are negligible and can be neglected here.:
[…]
where the covariant derivative $D_{\mu}=\partial_{\mu}-ig_{d}G^{d}_{\mu}$ is defined with $g_{d}$ and $G^{d}_{\mu}$ representing the $SU(3)_{D}$ gauge coupling and dark gluon field respectively. The field strength tensor $G^{d}_{\mu\nu}$ corresponds to the $SU(3)_{D}$ gauge field $G^{d}_{\mu}$, while $m_{q_{d}}$ and $m_{X}$ denote the dark quark and mediator masses. The coupling matrix $\kappa$ (a $n\times 3$ matrix) corresponds to interactions between $n$-flavor dark quarks and three-flavor charged leptons. Note that the mediator $X$ plays a crucial role in generating asymmetries in both SM and dark sectors through its decays, as it carries both SM lepton number and dark baryon number.
Figure 1: Diagrams illustrating dark shower (DS) production evolving into invisible or emerging lepton jets. Left panel: $e^{+}e^{-}$ annihilation via $t$-channel mediator $X$ producing a dark quark pair ($q_{d}\overline{q_{d}}$). Right panel: $q\overline{q}$ annihilation yielding mediator pairs, where each mediator decays to a dark quark and a SM lepton.
At lepton and hadron colliders, dark quarks are produced through $t$-channel processes and/or mediator decays, as shown in Fig. 1. Post-production, they acquire significant kinetic energy ($E\gg\Lambda_{d}$), where $\Lambda_{d}$ denotes the dark confinement scale. Assuming the dark QCD sector mirrors SM QCD dynamics, these dark quarks undergo dark parton shower evolution and hadronization, forming a collinear bunch of dark mesons.
In subsequent sections, we adopt the parameter settings for the QCD-like dark sector from Refs. Schwaller:2015gea ; Mies:2020mzw ; Archer-Smith:2021ntx ; Carrasco:2023loy to enable concrete simulation predictions. We assume three dark colors ($N_{C_{d}}=3$) and seven dark flavors ($N_{f_{d}}=7$), with the following mass relations: $\Lambda_{d}=m_{q_{d}}=2m_{\pi_{d}}$, and $m_{\rho_{d}}=4m_{\pi_{d}}$. Note that the dark confinement scale $\Lambda_{d}$ is closely related to the mass of the lightest dark baryon, which serves as the DM candidate Bai:2013xga ; Schwaller:2015gea ; Renner:2018fhh ; Mies:2020mzw ; Zhang:2021orr . All dark quarks are taken to be mass-degenerate, while dark pions and dark rho mesons are similarly degenerate in both flavor diagonal and off-diagonal states. Other dark baryons and mesons, assumed substantially heavier than dark pions and rho mesons, exhibit negligible production rates from dark showers and hadronization due to mass suppression. We therefore exclude them from this analysis.
Integrating out the heavy mediator (final terms in Eq. (1)) yields a dimension-six operator,
$\mathcal{L}_{\text{dim-6}}\supset\frac{\kappa^{2}}{m^{2}_{X}}(\overline{q_{d}}_{L}\ell_{R})(\overline{\ell_{R}}q_{dL}),$ (2)
which enables dark pion ($\pi_{d}$) decays to charged lepton pairs. The decay width is Schwaller:2015gea :
$\Gamma(\pi_{d}\to\bar{\ell}\ell)=\frac{\kappa^{4}f_{\pi_{d}}^{2}m_{\ell}^{2}m_{\pi_{d}}}{32\pi m_{X}^{4}},\quad\text{with}\quad\ell=e,\mu,\tau.$ (3)
Here $f_{\pi_{d}}$ is the dark pion decay constant and $m_{\ell}$ the lepton mass. Parameterizing the proper lifetime as
$c\tau_{0}=\frac{c\hbar}{\Gamma_{\pi_{d}}}\approx 1.92\,\text{m}\times\left(\frac{1}{\kappa^{4}}\right)\left(\frac{1\,\text{GeV}}{f_{\pi_{d}}}\right)^{2}\left(\frac{0.1\,\text{GeV}}{m_{\ell}}\right)^{2}\left(\frac{1\,\text{GeV}}{m_{\pi_{d}}}\right)\left(\frac{m_{X}}{1\,\text{TeV}}\right)^{4},$ (4)
reveals that dark pions propagate measurable distances from the primary vertex (PV) before decaying into charged lepton pairs.
Figure 2: The dark pion proper decay length with respect to $m_{\pi_{d}}$. Here we fixed the parameters $\kappa=1.0$, $f_{\pi_{d}}=m_{\pi_{d}}$, and $m_{X}=1$ TeV.
To illustrate the typical proper decay length of $\pi_{d}$ versus $m_{\pi_{d}}$, we fix $\kappa=1.0$, $f_{\pi_{d}}=m_{\pi_{d}}$, and $m_{X}=1$ TeV as benchmark values. In order to simplify our discussions here, the lepton flavor university (LFU) is assumed such that the common coupling $\kappa$ applies to the interactions of mediator with a dark quark and a SM charged lepton for each lepton flavor. Figure 2 shows that for $2m_{e}<m_{\pi_{d}}<2m_{\mu}$, $\pi_{d}$ decays to electron-positron pairs. The proper decay length exceeds $10^{7}$ m, allowing electrons and positrons to escape detection. When $2m_{\mu}<m_{\pi_{d}}<2m_{\tau}$, the muon decay channel opens, with $\pi_{d}$ predominantly decaying into muon pairs. The resulting decay length $\sim 1$ m makes these long-lived $\pi_{d}$ detectable. Finally, for $m_{\pi_{d}}>2m_{\tau}$, the tau channel becomes accessible, and $\pi_{d}$ primarily decays to tau pairs, potentially exhibiting prompt decays.
III Existing collider constraints on model parameters
Figure 2 reveals a dramatic change in the proper decay length of $\pi_{d}$ at the $m_{\pi_{d}}\geq 2m_{\mu}$ threshold. We therefore classify two dark pion mass ranges: $10\text{ MeV}\leq m_{\pi_{d}}<2m_{\mu}$ and $2m_{\mu}\leq m_{\pi_{d}}<2m_{\tau}$, and discuss their respective existing collider constraints.
Figure 3: Exclusion region for dark pion masses $10\text{ MeV}\leq m_{\pi_{d}}<2m_{\mu}$. The red dashed line separates long-lived (below) and prompt (above) $X^{\pm}$ decays. The blue and black dashed lines indicate where all $\pi_{d}$ particles escape detection at the LHC and LEP, respectively. The blue and black shaded regions correspond to slepton search constraints (LHC) ATLAS:2019lff and mono-photon search constraints (LEP) Fox:2011fx ; Liu:2019ogn . The green region represents the parameter space where the proper lifetime of $\pi_{d}$ exceeds 1 second, conservatively evaluated at $m_{\pi_{d}}=0.2\text{ GeV}$.
Figure 4: Exclusion region for dark pion masses $2m_{\mu}<m_{\pi_{d}}<2m_{\tau}$. The blue dashed line indicates where all $\pi_{d}$ particles escape detection at the LHC, making this region subject to slepton search constraints. The red dashed line retains the same interpretation as in Fig. 3. The green region represents the parameter space where the proper lifetime of $\pi_{d}$ exceeds 1 second, conservatively evaluated at $m_{\pi_{d}}=3\text{ GeV}$.
First, we consider the production of scalar mediator pairs through $pp\to X^{+}X^{-}$, with decays $X^{+}\to\ell^{+}q_{d}$ and $X^{-}\to\ell^{-}\overline{q_{d}}$ ($\ell=e,\mu,\tau$). Here, $\kappa$ in Eq. (1) affects only the $X^{\pm}$ lifetime, not the production cross-section. When $X^{\pm}$ decays promptly and dark parton showers are invisible at ATLAS and CMS detectors, signatures resemble slepton and stau searches in supersymmetry models ATLAS:2019lff ; ATLAS:2024fub . Under the LFU assumption, we recast these constraints onto this model. We generate a UFO model file for Eq. (1) via FeynRules Alloul:2013bka and simulate signal processes with MadGraph5_aMC@NLO Alwall:2014hca (version 3.5.8). For each $m_{X}$, we scan $\kappa$ and define the mediator prompt decay region by $\gamma\beta c\tau_{X}<0.1\text{ mm}$ in the lab frame. Figures 3 and 4 show regions above the red dashed line ($\kappa\gtrsim\mathcal{O}(10^{-7}-10^{-6})$) for the mediator prompt decays recast from slepton and stau constraints [^2] The signal production cross-sections rescaled to match Fig. 3 of Ref. Zhang:2021orr with QCD NLO+NLL corrections at $13\text{ TeV}$ LHC Fiaschi:2018xdm .. Additionally, because there are plenty of dark pions produced after dark showers and hadronization, we require the $\pi_{d}$ proper decay length (without any boost) larger than $10\text{ m}$ to conservatively assume all dark pions escape ATLAS and CMS detectors as missing energy. Specifically, we set $m_{\pi_{d}}=f_{\pi_{d}}=0.2\text{ GeV}$ in Fig. 3 and $m_{\pi_{d}}=f_{\pi_{d}}=3\text{ GeV}$ in Fig. 4 to estimate the upper bounds of $\kappa$. Regions below blue dashed lines ($\kappa\lesssim\mathcal{O}(1-10)$ for $10\text{ MeV}\leq m_{\pi_{d}}<2m_{\mu}$; $\kappa\lesssim\mathcal{O}(0.1-1)$ for $2m_{\mu}\leq m_{\pi_{d}}<2m_{\tau}$) correspond to invisible dark showers which are adapted for recasting.
Under the LFU assumption, the mediator $X^{\pm}$ has approximately equal decay branching ratios of $1/3$ to each flavor channel ($e^{\pm}q_{d}$, $\mu^{\pm}q_{d}$, $\tau^{\pm}q_{d}$). Consequently, we include an extra factor of $1/9$ when applying stau pair production constraints. Comparison with ATLAS results from Table 01 of auxiliary material in Ref. ATLAS:2024fub shows that for $m_{X}>200$ GeV and the dark quark mass $m_{q_{d}}=1$ GeV, the effective cross-section is at least a factor of two below the experimental bounds. Thus, no exclusion region exists for $m_{X}>200$ GeV, though stau constraints exclude $m_{X}\lesssim 160$ GeV.
For slepton pair production constraints in Fig. 3c of Ref. ATLAS:2019lff , the analysis involves all combinations of a pair of leptons from electrons and muons in the final states. Therefore, we apply a factor of $4/9$ in our process. Notably, while the constraints include contributions from both $\tilde{\ell}^{+}_{L}\tilde{\ell}^{-}_{L}$ and $\tilde{\ell}^{+}_{R}\tilde{\ell}^{-}_{R}$ production, our signal process specifically corresponds to the $\tilde{\ell}^{+}_{R}\tilde{\ell}^{-}_{R}$ channel. We therefore isolate the right-handed slepton constraints using the fraction of $pp\to\tilde{\ell}^{+}_{L}\tilde{\ell}^{-}_{L}$ and $pp\to\tilde{\ell}^{+}_{R}\tilde{\ell}^{-}_{R}$ production cross-sections from Refs. Fuks:2013lya ; Fiaschi:2018xdm . The resulting effective cross-section in our process exceeds ATLAS exclusion limits for $\tilde{\ell}^{+}_{R}\tilde{\ell}^{-}_{R}$ in the mass range $m_{X}\in[200,700]$ GeV. Additionally, after accounting for the average Lorentz boost of the mediator ($\gamma\beta\simeq 1.5$), we estimate the prompt decay boundaries within this mass range.
Conversely, for $\pi_{d}$ with moderate decay lengths, the novel signatures of displaced lepton jets are expected to be observable at the LHC. However, Ref. Zhang:2021orr notes detection challenges for the $pp\to X^{+}X^{-}$ production channel due to: (1) insufficient production cross-sections, and (2) detection efficiencies after considering the whole cut-flow below $\sim 0.1$. Moreover, significant irreducible cosmic-ray backgrounds further complicate the observation. While LHC Run 3 and HL-LHC may improve sensitivity, current ATLAS displaced lepton jet searches ATLAS:2022izj exclusively target Higgs exotic decays, distinct from this model’s signature of two opposite-sign leptons plus two displaced lepton jets. Consequently, recasting strategies from Ref. ATLAS:2022izj are inapplicable. A dedicated analysis (beyond the scope of this work) would be required, though we establish that the LHC constraints remain weaker than those derived for $e^{+}e^{-}$ colliders in Sections IV and V.
Dark quark pairs can also be produced via $t$-channel mediator exchange at the LEP. Following dark showers and hadronization, if all dark mesons escape the detector, they contribute to missing energy. With initial state radiation (ISR), this also yields a mono-photon signature. For $10\text{ MeV}\leq m_{\pi_{d}}<2m_{\mu}$, we conservatively set $m_{\pi_{d}}=f_{\pi_{d}}=0.2\text{ GeV}$ and require $\pi_{d}$ proper decay length greater than $12$ m to ensure that all dark mesons escape the detector volume of LEP. The black dashed line in Fig. 3 gives upper bounds $\kappa\sim\mathcal{O}(1-10)$. We further recast mono-photon constraints at the LEP from the $t$-channel scalar mediator and Dirac fermionic DM model with the cut-off scale $\Lambda$ (Eq. 4) and the purple line in Fig. 4 ($\bar{\chi}e\bar{e}\chi$) of Ref. Fox:2011fx to the leptophilic composite ADM via:
$\frac{1}{\Lambda^{2}}\simeq\frac{1}{(333.86\text{ GeV})^{2}}=3\frac{\kappa^{2}}{m^{2}_{X}},$ (5)
where the factor of 3 originates from $SU(3)_{D}$ charges. Assuming light dark quarks and determining lower bounds of $\Lambda$, the black regions in Fig. 3 indicate where predicted cross-sections exceed experimental bounds, consistent with the $1/\Lambda^{2}$ scaling.
Finally, a dark pion lifetime exceeding one second would violate observed light-element abundances and the CMB spectrum, as constrained by Big Bang Nucleosynthesis (BBN) and cosmic microwave background (CMB) measurements Lindley:1984bg ; Dev:2025pru . We therefore incorporate these constraints by excluding regions where the proper lifetime of $\pi_{d}$ is greater than 1 sec, indicated by the green shaded areas in Figs. 3 and 4. Notably, only a narrow region of parameter space remains in Fig. 3 for $m_{X}\lesssim 1$ TeV.
Although portions of the parameter space are excluded in Figs. 3 and 4, significant regions remain accessible for exploration at both low-energy and high-energy $e^{+}e^{-}$ colliders. In subsequent sections, we derive exclusion limits for three benchmark points (BPs): $m_{\pi_{d}}=0.32$, $0.7$, and 11 GeV.
IV Sparse dark showers at BaBar, Belle II and GAZELLE
As previously discussed, the intensity of dark showers depends critically on the energy of the produced dark quarks. This section focuses on low-energy $e^{+}e^{-}$ colliders capable of generating sparse dark showers, where individual displaced vertices from long-lived dark pion decays can be resolved. We specifically examine the BaBar and Belle II experiments, along with the proposed GAZELLE far detector for Belle II.
In this and the following section, we employ FeynRules Alloul:2013bka to generate a UFO model file for Eq. (1), and then parton-level Monte Carlo events are generated using MadGraph5_aMC@NLO Alwall:2014hca . Subsequently, the Hidden Valley module within Pythia8.313 Sjostrand:2014zea simulates dark parton showering and hadronization processes. Parameter settings closely follow Refs. Schwaller:2015gea ; Mies:2020mzw ; Archer-Smith:2021ntx ; Carrasco:2023loy , specifically: $N_{C_{d}}=3$, $N_{f_{d}}=7$, $\Lambda_{d}=m_{q_{d}}=2m_{\pi_{d}}$, and $m_{\rho_{d}}=4m_{\pi_{d}}$. In our simulation, diagonal and off-diagonal HV-mesons share identical decay patterns. We implement the running coupling $\alpha_{d}$ and apply a $p_{T}$ cutoff of $1.1\Lambda_{d}$ to dark showers. The fraction of vector and pseudoscalar HV-mesons is calculated by following Appendix A of Ref. Knapen:2021eip . Finally, the ratio of dark pion mass over dark pion decay constant, $m_{\pi_{d}}/f_{\pi_{d}}$, can range from $\mathcal{O}(1-10)$ Berlin:2018tvf ; Choi:2018iit ; Kuwahara:2023vfc ; Bernreuther:2023kcg . For this analysis, we fix $f_{\pi_{d}}=m_{\pi_{d}}$. This choice only affects the dark pion decay width, which can be easily rescaled using Eq. (3).
IV.1 Recasting LLP searches at the BaBar
The BaBar experiment utilized the PEP-II collider, colliding $9\,\text{GeV}$ electrons with $3.1\,\text{GeV}$ positrons at a center-of-mass energy $\sqrt{s}=10.58\,\text{GeV}$. Dark quark pairs are produced via the $t$-channel process (Fig. 1, left panel). Following dark showering and hadronization, long-lived dark pions decay to muon pairs, creating multiple displaced vertices. Such novel signature is an icon of dark showers.
To recast the previous BaBar long-lived particle (LLP) searches to this model, we implement the event selection criteria from Refs. BaBar:2015jvu ; Bernreuther:2022jlj :
  • The transverse displacement of $\pi_{d}$ is required to be within the range 10 mm $R_{xy}<$ 500 mm from the interaction point (IP).
  • The final state must include at least three charged tracks with $p_{T}>0.12$ GeV, among which at least two of them must have $p_{T}>0.18$ GeV.
  • The invariant mass of the decay products must satisfy either $m_{\mu^{+}\mu^{-}}<0.37$ GeV or $m_{\mu^{+}\mu^{-}}>0.5$ GeV.
  • Each muon is required to have $d_{0}/\sigma_{d_{0}}>3$, where $d_{0}$ is the transverse impact parameter and $\sigma_{d_{0}}$ is its uncertainty.
The detector resolution is performed by the smearing effect of the relevant uncertainties of observables. Since recasting efficiencies are only available for LLPs with $m>0.5~\text{GeV}$ and $c\tau>5~\text{mm}$ as noted in Refs. BaBar:2015jvu ; Bernreuther:2022jlj , we restrict analysis to two of our BPs: $m_{\pi_{d}}=0.7~\text{GeV}$ and $1~\text{GeV}$, requiring $c\tau>5~\text{mm}$ for dark pions.
Specifically, to calculate the impact parameter $d_{0}$, we first select the long-lived dark pion candidate and its daughter muons. We record their respective three-momenta and compute the sine of the angle $\alpha$ between them. The transverse displacement is defined as $R_{xy}=\sqrt{R_{x}^{2}+R_{y}^{2}}$, where $R$ is the flight distance of the dark pion in the lab frame, and $R_{x}$ ($R_{y}$) is the $x$ ($y$) component of $R$. The impact parameter is then given by $d_{0}=R_{xy}\sin\alpha$. The position resolution of displaced vertices and the momentum resolution of the final-state particles are given by BaBar:1995bns
$\sigma_{xy}=\sigma_{z}=\left[\frac{50\text{ GeV}}{p_{T}}+15\right]\mu\text{m},$ (6)
$\frac{\sigma_{p_{T}}}{p_{T}}=\left[0.21+0.14\frac{p_{T}}{\text{GeV}}\right]\%.$ (7)
Here, $p_{T}$ denotes the transverse momentum of the charged track. In the calculation of $d_{0}$, we use $\sigma_{xy}$ to account for the detector resolution in the transverse plane.
Figure 5: Kinematic distributions for the signal benchmark points BP2 and BP3 at the BaBar. The distributions of the leading $p_{T}$ and subleading $p_{T}$ dark pions and their daughter muons are represented by solid and dashed lines, respectively. Top-left: transverse momentum $p_{T}$ of muons from dark pion decays; Top-right: invariant mass $m_{\mu^{+}\mu^{-}}$ of muon pairs; Bottom-left: transverse decay distance $R_{xy}$ of dark pions; Bottom-right: track impact parameter significance $d_{0}/\sigma_{0}$.
Figure 5 shows the relevant kinematic distributions of these two signal BPs. First, the transverse momentum distribution of the muon, $P_{T}(\mu)$ (top-left), is generally larger for heavier $\pi_{d}$ masses. This is because, under a lower total $\pi_{d}$ yield, each $\pi_{d}$ carries higher average energy, leading to muons from its decay with larger $p_{T}$. Second, in the $\mu^{+}\mu^{-}$ invariant mass distribution, $m_{\mu^{+}\mu^{-}}$ (top-right), the position of the reconstructed mass peak directly reflects the mass of $\pi_{d}$. Regarding the transverse decay length, $R_{xy}$ (bottom-left), since $\pi_{d}$ is long-lived, its decay vertex is significantly displaced from the IP in the transverse plane. Moreover, due to the proper decay length $c\tau_{0}\propto 1/m_{\pi_{d}}^{3}$, a lighter $\pi_{d}$ has a longer decay length, resulting in an $R_{xy}$ distribution that extends to larger values and exhibits a more pronounced long tail for BP2 compared to BP3. Finally, the transverse impact parameter $d_{0}$ (bottom-right) further characterizes the long-lived nature of $\pi_{d}$. The value of $d_{0}$ is significantly larger than the uncertainty $\sigma_{0}$ which clearly distinguishes $\pi_{d}$ decays from prompt decays.
Figure 6: Exclusion limits for three benchmark points, $m_{\pi_{d}}=0.32$, $0.7$ and $1$ GeV for three experiments: BaBar, Belle II and L-GAZELLE.
The production cross-section for $e^{+}e^{-}\to\overline{q_{d}}q_{d}$ scales as $(\kappa/m_{X})^{4}$, while the proper lifetime of $\pi_{d}$ is inversely proportional to this factor. We therefore recast BaBar LLP search constraints onto the $(m_{X},\kappa)$ parameter plane. Specifically, we extract exclusion limits from the $\mu^{+}\mu^{-}$ channel (red curve in Fig. 2 of Ref. BaBar:2015jvu ), plotting LLP mass versus $\sigma(e^{+}e^{-}\to LX)\mathcal{B}(L\to f)\epsilon(f)$, where $\sigma(e^{+}e^{-}\to LX)$ is the LLP production cross-section ($L$ is the LLP candidate, and $X$ are associated particles); $\mathcal{B}(L\to f)$ is the branching fraction to final state of a pair of $f$ particles (a muon pair in our case); $\epsilon(f)$ is the detection efficiency to such final state. For $m_{\pi_{d}}=0.7$ GeV and $1$ GeV, exclusion limits appear in gray regions of Fig. 6. The upper boundaries correspond to short dark pion lifetimes that preclude LLP classification, while lower boundaries indicate long-lived pions escaping the detector’s sensitive volume. The inverse relationship of constraints between $m_{X}$ and $\kappa$ is clearly evident in Fig. 6.
IV.2 Detecting dark showers at the Belle II
The Belle II experiment operates at the SuperKEKB collider, which collides $7\text{ GeV}$ electrons with $4\text{ GeV}$ positrons at a center-of-mass energy of $\sqrt{s}=10.58\text{ GeV}$. Dark quark pairs are produced via the process $e^{+}e^{-}\to\overline{q_{d}}q_{d}$, followed by dark showers and hadronization, as illustrated in the left panel of Fig. 1. This leads to the production of several long-lived dark pions, which predominantly decay into muon pairs. As a result, multiple displaced vertices can be identified as the novel signal signature. Three BPs with $m_{\pi_{d}}=0.32$, $0.7$, and $1$ GeV are chosen for illustration, corresponding to $\kappa=1.0$ and $m_{X}=200\text{ GeV}$.
$m_{\pi_{d}}$ [GeV] $\langle{N}_{\pi_{d}}\rangle$ $\langle{E}(\pi_{d})\rangle$ [GeV] $\langle{P_{T}}(\mu)\rangle$ [GeV]
0.32 6.20 1.78 0.63
0.7 3.15 3.46 1.26
1 2.03 5.43 1.96
Table 1: The average dark pion $\pi_{d}$ multiplicity, the average energy of each $\pi_{d}$, and the average transverse momentum of each muon in the final state for three BPs with $m_{\pi_{d}}=0.32$, $0.7$, and $1$ GeV at the Belle II.
Trigger Conditions
Two tracks Two tracks with $p_{T}>0.3$ GeV for each and an azimuthal opening angle $\Delta\phi>90^{\circ}$ at the interaction point, as well as the angle between the position vector of decay vertex and the beam axis with $38^{\circ}<\theta_{\text{lab}}<127^{\circ}$ in the lab frame
One muon One muon with $p_{T}>0.9$ GeV and $38^{\circ}<\theta_{\text{lab}}<127^{\circ}$
Displaced vertex At least one displaced vertex in the event with $9$ mm <R<600<R<600 mm, formed from two tracks with $p_{T}>0.1$ GeV and $38^{\circ}<\theta_{\text{lab}}<127^{\circ}$
Table 2: Three Belle II triggers are considered in this analysis: two tracks, one muon, and displaced vertex, adapted from Ref. Bernreuther:2022jlj .
Object Selection Criteria
Muons $p_{T}(\mu^{+}),p_{T}(\mu^{-})>0.05$ GeV $m_{\mu^{+}\mu^{-}}<0.48$ GeV or $m_{\mu^{+}\mu^{-}}>0.52$ GeV
Displaced vertices $0.2$ cm $R_{xy}<60$ cm 55-55 cm $\leq z\leq 140$ cm $17^{\circ}\leq\theta_{\text{lab}}\leq 150^{\circ}$
Table 3: The event selection criteria for muons and displaced vertices in LLP searches at the Belle II Bernreuther:2022jlj .
In all BPs, the average number of $\pi_{d}$ per event exceeds two as shown in Table 1. Thus, we require at least two $\pi_{d}$’s per event for the signal process in subsequent analyses. All events must satisfy the Belle II trigger selections outlined in Table 2 and then pass the event selection criteria outlined in Table 3 Bernreuther:2022jlj . Moreover, key kinematic properties for each BP are summarized in Table 1, including the average $\pi_{d}$ multiplicity, the average energy of each $\pi_{d}$, and the average transverse momentum of each muon in the final state.
Trigger Type $m_{\pi_{d}}=0.32\,\text{GeV}$ $m_{\pi_{d}}=0.70\,\text{GeV}$ $m_{\pi_{d}}=1.00\,\text{GeV}$
Two Tracks 75.8275.82% 58.7258.72% 34.1634.16%
After Cuts 25.0425.04% 47.4847.48% 25.3125.31%
One Muon 83.7683.76% 89.2289.22% 84.1784.17%
After Cuts 13.7413.74% 65.0165.01% 51.6251.62%
Displaced Vertex 93.5693.56% 79.4179.41% 46.2446.24%
After Cuts 15.5615.56% 60.1260.12% 35.1635.16%
Table 4: Trigger efficiencies for three different dark pion masses with fixed $\kappa=1.0\,,m_{X}=200\,\text{GeV}$.
Figure 7: Distributions for leading (solid line) and subleading (dashed line) $p_{T}$ dark pions ($\pi_{d}$) in BP1, BP2, and BP3 at the Belle II. Top-left: Number of dark pions per event, $N_{\pi_{d}}$; Top-right: Transverse decay length, $R_{xy}$, of leading and subleading $\pi_{d}$’s; Bottom-left: Longitudinal flight distance, $R_{z}$; Bottom-right: Angle between the decay vertex vector and beam axis, $\theta_{\text{lab}}$.
In this study, the trigger strategy at the Belle II relies on three categories as follows:
  • Two tracks trigger: An event passes this trigger if it contains two muon tracks, each with $p_{T}>0.3\text{ GeV}$, and the azimuthal opening angle between them at the IP in the lab frame satisfies $\Delta\phi>90^{\circ}$. In addition, the angle between the position vector of decay vertex and the beam axis in the lab frame must satisfy $38^{\circ}<\theta_{\text{lab}}<127^{\circ}$.
  • One muon trigger: An event passes this trigger if it contains at least one muon with $p_{T}>0.9\text{ GeV}$, and $38^{\circ}<\theta_{\text{lab}}<127^{\circ}$.
  • Displaced vertex trigger: This trigger is satisfied when the event contains at least one LLP (dark pion) with a transverse decay length $9~\text{mm}<R_{xy}<600~\text{mm}$, which decays into two muons, each with $p_{T}>0.1\text{ GeV}$, and $38^{\circ}<\theta_{\text{lab}}<127^{\circ}$.
Figure 8: The kinematic distributions for muons from leading $p_{T}$ (solid line) and subleading $p_{T}$ (dashed line) dark pions in BP1, BP2, and BP3 at the Belle II. Top-left: Transverse momentum of muons, $p_{T}(\mu)$; Top-right: Invariant mass of muon pairs, $m_{\mu^{+}\mu^{-}}$; Bottom-left: Energy of muon pairs, $E_{\mu^{+}\mu^{-}}$; Bottom-right: Angular separation between opposite-sign muons, $\Delta R_{\mu^{+}\mu^{-}}$.
We further apply the event selection criteria summarized in Table 3 to identify the signal events and suppress the background ones. An event is selected if at least two leading-$p_{T}$ $\pi_{d}$ particles satisfy the following conditions: $0.2~\text{cm}<R_{xy}<60~\text{cm}$, $55~\text{cm}\leq z\leq 140~\text{cm}$, and $17^{\circ}\leq\theta_{\text{lab}}\leq 150^{\circ}$. Additionally, their daughter muons must satisfy $p_{T}>0.05~\text{GeV}$, and the reconstructed invariant mass of the muon pair must fulfill $m_{\mu^{+}\mu^{-}}<0.48~\text{GeV}$ or $m_{\mu^{+}\mu^{-}}>0.52~\text{GeV}$.
Table 4 summarizes the efficiencies of three trigger types and after applying event selection criteria for three BPs. Overall, the efficiencies of all three triggers exhibit significant variations with $m_{\pi_{d}}$, and the relative ordering of efficiencies among the BPs changes considerably after the same selection is applied. Specifically, the efficiency of Two tracks trigger decreases as $m_{\pi_{d}}$ increases, dropping from $75.82\%$ for BP1 to $34.16\%$ for BP3. This behavior can be understood from the distribution of $\theta_{\text{lab}}$. For $m_{\pi_{d}}=1~\text{GeV}$, the tracks from the leading and sub-leading $p_{T}$ $\pi_{d}$’s are more dispersed in $\theta_{\text{lab}}$ distribution, with many falling outside the trigger acceptance region of $38^{\circ}\leq\theta_{\text{lab}}\leq 127^{\circ}$. In this case, the two selected tracks are very likely to originate from the same dark pion, and their opening angle $\Delta\phi$ rarely exceeds $\pi/2$, leading to the lowest efficiency. In contrast, for $m_{\pi_{d}}=0.32$ and $0.7~\text{GeV}$, the $\theta_{\text{lab}}$ distribution is more concentrated, and the higher $\pi_{d}$ multiplicity results in higher efficiency. One muon trigger maintains high efficiency across all three BPs, peaking at $89.22\%$ for BP2. The efficiency of Displaced vertex trigger decreases with increasing $m_{\pi_{d}}$. This condition strongly depends on the decay length of $\pi_{d}$. The proper decay lengths for the three BPs are $c\tau_{0}=87.87~\text{mm}$ (BP1), $8.40~\text{mm}$ (BP2), and $2.88~\text{mm}$ (BP3), respectively. The proper decay length of BP1 is an order of magnitude larger than that of BP2, after multiplying by $\beta\gamma$ in the lab frame, it often exceeds the upper limit of the trigger condition. However, since the proper decay length follows an exponential distribution in the Monte Carlo simulation, the high multiplicity of BP1 enhances the probability that at least one displaced vertex satisfies the trigger condition. In contrast, for $m_{\pi_{d}}=1~\text{GeV}$, the $R_{xy}$ distribution of the two leading $p_{T}$ $\pi_{d}$’s is concentrated in a very small region, resulting in the fewest events satisfying the trigger condition. After applying the cuts in Table 3, the order of efficiencies in these three cases may change. From the $R_{xy}$ distribution of $\pi_{d}$’s, for $m_{\pi_{d}}=0.32~\text{GeV}$, many events have decay lengths for both $\pi_{d}$’s exceeding 60 cm, leading to a significant reduction in efficiency. In comparison, for $m_{\pi_{d}}=0.7~\text{GeV}$, most events fall within the required range, resulting in the highest efficiency.
We then analyze relevant signal kinematic distributions for $\pi_{d}$’s and muons at the Belle II in Figs. 7 and 8. First, as the $m_{\pi_{d}}$ increases from BP1 to BP3, the $N_{\pi_{d}}$ distribution (top-left) decreases from approximately 6 to 2. Regarding the decay length, $R_{xy}$ and $R_{z}$ represent the transverse and longitudinal decay lengths of the dark pions in the lab frame, respectively. According to Eq. (4), under fixed parameters and with $f_{\pi_{d}}=m_{\pi_{d}}$, the proper decay length satisfies $c\tau_{0}\propto 1/m_{\pi_{d}}^{3}$. Therefore, lighter $\pi_{d}$’s have longer decay lengths, and their $R_{xy}$ (top-right) and $R_{z}$ (bottom-left) distributions exhibit larger values and longer tails. Finally, the $\theta_{\rm lab}(\text{track})$ distribution (bottom-right) is significantly dispersed for the BP1, while the other two BPs exhibit relatively concentrated distributions. By comparing $E$ and $p_{T}$ of daughter particles, it can be seen that for BP1 and BP2, the leading and subleading $p_{T}$ $\pi_{d}$’s have similar $p_{z}$ values, leading to small differences in their $\Delta\theta_{\rm lab}$ distributions. In contrast, for the BP3, although its $p_{T}$ distribution is similar, the leading $\pi_{d}$ carries higher energy, resulting in a larger $p_{z}$ difference. The greater difference in boost further leads to a significant separation in their $\theta_{\rm lab}$ distributions.
As shown in Fig. 8, the $p_{T}(\mu)$ distribution (top-left) is similar to that in the BaBar analysis. In the $m_{\mu^{+}\mu^{-}}$ distribution (top-right), the reconstructed mass peaks clearly appear near the respective $m_{\pi_{d}}$: $0.32$ GeV, $0.7$ GeV, and $1$ GeV for BP1, BP2, and BP3, respectively, but exhibit different levels of broadening. The width of these distributions is influenced by momentum resolutions of muons Adachi:2018qme ; Kang:2021oes :
$\frac{\sigma_{p}}{p_{\mu}}=0.0011p_{\mu}[\text{GeV}]\oplus\frac{0.0025}{\beta},$ (8)
the width of the $m_{\mu^{+}\mu^{-}}$ distribution increases from BP1 to BP3 as the momentum resolution grows with the $p_{T}(\mu)$. The $E_{\mu^{+}\mu^{-}}$ distribution (bottom-left) shows a clear $m_{\pi_{d}}$ and $p_{T}(\mu)$ dependence: muon pairs from the decay of heavier and higher-$p_{T}$ $\pi_{d}$’s carry more energy. Finally, in the $\Delta R_{\mu^{+}\mu^{-}}$ distribution (bottom-right), the average angular separation between the muons is smaller for higher $\pi_{d}$ multiplicity, while in BP3, which has fewer $\pi_{d}$’s, the $\Delta R$ distribution is generally larger.
Since the coupling $\kappa$ and mediator mass $m_{X}$ affect both the $e^{+}e^{-}\to\overline{q_{d}}q_{d}$ production cross-section and $\pi_{d}$ lifetime, we map exclusion limits in the parameter space of $(m_{X},\kappa)$ plane. In particular, $\sigma(e^{+}e^{-}\to\overline{q_{d}}q_{d})$ and $\Gamma_{\pi_{d}}$ both scale as $(\kappa/m_{X})^{4}$. Applying the trigger conditions and event selection criteria (Tables 2 and 3), we obtain the signal efficiency $\epsilon_{S}$. Here an event is considered to pass the trigger selection if at least one of the three trigger conditions is satisfied. The expected signal yield is $N_{S}=\mathcal{L}\times\sigma(e^{+}e^{-}\to\overline{q_{d}}q_{d})\times\epsilon_{S}$, assuming optimistic integrated luminosity $\mathcal{L}=50~\text{ab}^{-1}$ for the Belle II experiment. With a background-free assumption, $N_{S}=2.3$ events establish the exclusion contours of the $90\%$ confidence level (CL) which correspond to the lower boundaries with purple solid lines in Fig. 6.
As expected, small $\kappa/m_{X}$ yields a tiny cross-section and excessive $\pi_{d}$ decay length, reducing $N_{S}$ below detection ability of the Belle II experiment. Conversely, large $\kappa/m_{X}$ produces prompt $\pi_{d}$ decays without the signature of displaced vertices. In this situation, the efficiency drops sharply, necessitating enormous Monte Carlo samples. Therefore, we conservatively set $N_{S}=10$ to estimate upper boundaries with purple dashed lines in Fig. 6333Even with this relaxed requirement, more than $10^{8}$ simulated events are required for this estimation.. We can observe that the projected Belle II constraints also indicate inverse proportionality between $m_{X}$ and $\kappa$ for these three BPs. Notably, blank regions in the upper-left corner of $(m_{X},\kappa)$ plane could be covered by searching for multi-muons from the prompt decay of dark pions. However, detailed signal-background analysis is beyond the scope of this work.
Unlike most light DM searches at the Belle II, the mono-photon search strategy lacks sufficient discriminating power between signal and background events in this model. Although undetectable dark pions produce missing energy signatures, their kinematic distributions remain similar to SM backgrounds. Because the mediator mass exceeds the center-of-mass energy at the Belle II, this absence of light mediator production results in smooth, continuous photon energy distributions that resemble SM backgrounds Bernreuther:2022jlj . We therefore exclude mono-photon constraints from Belle II in this analysis.
IV.3 Long-lived dark pion searches in proposed GAZELLE detector
Positioned downstream of the Belle II detector at a considerable distance from the IP, the proposed GAZELLE detector facilitates long-distance tracking and displaced vertex reconstruction which enhances the overall sensitivity for probing LLPs. The GAZELLE detector has three proposed configurations: Baby-GAZELLE, L-GAZELLE, and GODZILLA Ko:2025drr . L-GAZELLE comprises two components, LG-B1 (Region I) and LG-B2 (Region II), which select signal events as detailed later. We chose L-GAZELLE for this analysis due to its superior angular coverage (LG-B1: 0.34 sr [2.7%]; LG-B2: 0.76 sr [6.0%]), exceeding Baby-GAZELLE (0.12 sr, 0.95%) and GODZILLA (0.74 sr, 5. 9%) and thereby improving the detection efficiency. Its moderate detection distances (LG-B1: $x=35$ m, $y=2.3$ m, $z=0$ m; LG-B2: $x=19$ m, $y=2.3$ m, $z=10.5$ m) optimally capture long-lived dark pion decays. This contrasts with Baby-GAZELLE’s proximity ($x=10$ m, y=3.7y=-3.7 m, $z=10$ m), which limits the far decay sensitivity, and the GODZILLA distance (x=27x=-27 m, $y=18$ m, $z=20$ m), which reduces the near decay sensitivity. Thus, L-GAZELLE offers the optimal balance for our study.
We apply Displaced Decay Counter (DDC) Domingo:2023dew for the L-GAZELLE detector simulation [^4] The L-GAZELLE detector is approximated with LG-B1 (single-layer approximation within $8.5\%$ error) and LG-B2 (four-layer approximation within $5.0\%$ error).. The decay probability of LLPs within a certain range is calculated as: $P(\ell_{1},\ell_{2})=\exp\left(-\frac{\ell_{1}}{\gamma|\vec{v}|\tau_{0}}\right)-\exp\left(-\frac{\ell_{2}}{\gamma|\vec{v}|\tau_{0}}\right)$, where $\gamma\equiv(1-|\vec{v}|^{2})^{-1/2}$, $|\vec{v}|$ and $\tau_{0}$ are the velocity and proper decay length of a LLP with the natural units. The weight for each layer is $\Delta\phi/(2\pi)$, where $\Delta\phi$ is the angular aperture of each layer. The weighted sum of probabilities from all detector layers gives the acceptance $A_{i}$ for each LLPi, which represents the detection probability of each LLP. For event selections, we calculate the probabilities of detecting at least one LLP and at least two LLPs:
  • At least one LLP detected: $\langle P(\geq 1)\rangle=1-\langle\prod(1-A_{i})\rangle$;
  • At least two LLPs detected: $\langle P(\geq 2)\rangle=1-\langle\prod(1-A_{i})\rangle-\langle\sum A_{i}\prod_{j\neq i}(1-A_{j})\rangle$.
For the signal efficiency $\epsilon_{S}$ obtained for both scenarios, the expected signal yield is given by $N_{S}=\mathcal{L}\times\sigma(e^{+}e^{-}\to\bar{q}_{d}q_{d})\times\epsilon_{S}$, assuming an optimistic integrated luminosity $\mathcal{L}=50~\text{ab}^{-1}$. The $90\%$ C.L. exclusion contours correspond to the lower boundaries, shown as orange-red (at least one LLP detected) and brown (at least two LLPs detected) solid lines in Fig. 6. Compared to recast BaBar and projected Belle II constraints, the far-detector L-GAZELLE coverage extends to smaller $\kappa$ and larger $m_{X}$ values.
$m_{\pi_{d}}$ [GeV] $\tau_{0}$ [m] $\langle P(\geq 1)\rangle$ $\langle P(\geq 2)\rangle$
0.32 54.91 $9.02\times 10^{-3}$ $3.71\times 10^{-5}$
10.98 $2.16\times 10^{-2}$ $1.96\times 10^{-4}$
5.49 $2.27\times 10^{-2}$ $1.99\times 10^{-4}$
0.70 5.25 $1.22\times 10^{-2}$ $2.91\times 10^{-5}$
1.05 $2.68\times 10^{-3}$ $8.37\times 10^{-8}$
0.53 $17\times 10^{-4} 3.3$$33\times 10^{-11}$
1.00 1.80 $4.90\times 10^{-3}$ $7.36\times 10^{-8}$
0.36 $7.32\times 10^{-6}$ $9.96\times 10^{-19}$
0.18 $2.02\times 10^{-9}$ 0
Table 5: Detection efficiency $\langle P(\geq 1)\rangle$ and $\langle P(\geq 2)\rangle$ for various dark pion proper decay length $\tau_{0}$ and dark pion mass $m_{\pi_{d}}$ values. Corresponding production cross-sections are $7.15\times 10^{-2}$ fb for $m_{\pi_{d}}=0.32$ GeV at $\tau_{0}=54.91$ m, $6.85\times 10^{-2}$ fb for $m_{\pi_{d}}=0.7$ GeV at $\tau_{0}=5.25$ m, and $6.45\times 10^{-2}$ fb for $m_{\pi_{d}}=1$ GeV at $\tau_{0}=1.80$ m.
We further evaluated detection efficiencies for different $m_{\pi_{d}}$ and $\tau_{0}$ values, as shown in Table 5. The detection efficiency for at least two LLPs ($\langle P(\geq 2)\rangle$) is significantly lower than for at least one LLP ($\langle P(\geq 1)\rangle$). For BP1, $\langle P(\geq 2)\rangle$ is about two orders of magnitude smaller than $\langle P(\geq 1)\rangle$. As the $m_{\pi_{d}}$ increases, the efficiency gap widens, particularly at $m_{\pi_{d}}=1$ GeV, where events typically produce two LLPs that travel nearly back-to-back, making it highly unlikely for both to be detected within L-GAZELLE. Since $\tau_{0}$ and the production cross-section are anti-correlated, a decrease in $\tau_{0}$ enhances the cross-section but causes a more rapid decline in detection efficiency. This efficiency loss prevents meaningful constraints for BP2 and BP3 in the scenario requiring at least two detected LLPs. Consequently, the brown solid line corresponding to BP2 and BP3 is absent in Fig. 6.
V Dense dark showers at the CEPC
This section further investigates the detection prospects at a high-energy $e^{+}e^{-}$ collider, focusing on the proposed circular electron positron collider (CEPC) with a center-of-mass energy of $\sqrt{s}=240~\text{GeV}$ and an integrated luminosity of $5.6~\text{ab}^{-1}$. We also study the exclusion limits for two mass ranges of the dark pion: $10~\text{MeV}\leq m_{\pi_{d}}<2m_{\mu}$, and $2m_{\mu}\leq m_{\pi_{d}}<2m_{\tau}$, which lead to the mono-photon and displaced muon jet (DMJ) signatures.
V.1 The mono-photon signature
Figure 9: Exclusion regions for different dark pion mass ranges. The first panel shows the mono-photon signature for $10~\text{MeV}<m_{\pi_{d}}<2m_{\mu}$, while the remaining three panels show benchmark points $m_{\pi_{d}}=0.32$, $0.7$ and $1~\text{GeV}$ for both CEPC mono-photon and DMJ signatures, covering $2m_{\mu}<m_{\pi_{d}}<2m_{\tau}$.
For the dark pion mass range $10~\text{MeV}\leq m_{\pi_{d}}<2m_{\mu}$, we maintain the conservative setting $m_{\pi_{d}}=f_{\pi_{d}}=0.2~\text{GeV}$ and impose the requirement that the $\pi_{d}$ proper decay length exceeds $6.08~\text{m}$ to guarantee that all dark mesons decay outside the CEPC detector volume. The resulting upper bounds on $\kappa$, shown as the yellow dashed curve in Fig. 9, reach values of $\mathcal{O}(1-10)$. We recast the previous CEPC mono-photon analysis, derived for the $t$-channel scalar mediator effective operator with cutoff scale $\Lambda_{t}$ (corresponding to the $t$-channel case in Eq. (2.3)) and the upper right panel (H-mode) of Fig. 16 in Ref. Liu:2019ogn , to our leptophilic composite ADM framework through the relation:
$\frac{1}{\Lambda^{2}}\simeq\frac{1}{(1078~\text{GeV})^{2}}=3\frac{\kappa^{2}}{m^{2}_{X}}.$ (9)
Assuming light dark quarks and determining the lower limits on $\Lambda$, the yellow region in Fig. 9 indicates parameter space where the predicted cross-sections exceed the experimental bounds.
V.2 The displaced muon jet signature
For the dark pion mass range $2m_{\mu}\leq m_{\pi_{d}}<2m_{\tau}$, unlike low-energy $e^{+}e^{-}$ colliders, due to the high center-of-mass energy, the initially produced dark quarks undergo dark showering and hadronization processes, yielding a large number of dark pions with the same model parameter settings. These dark pions subsequently decay to muons, resulting in a high multiplicity of displaced muons that collectively form displaced muon jets (DMJs).
Detector Component $R_{\text{in}}$ $R_{\text{out}}$ $\sigma_{xy}$
[…]
Hadron Calorimeter $2.30\text{\,}\mathrm{m}$ $3.34\text{\,}\mathrm{m}$ $30.00\text{\,}\mathrm{m}\mathrm{m}$
Muon System $4.40\text{\,}\mathrm{m}$ $6.08\text{\,}\mathrm{m}$ $20.00\text{\,}\mathrm{m}\mathrm{m}$
Table 6: The CEPC detector geometry and its relevant transverse spatial resolution Zhang:2021orr .
The detailed definition of DMJ in this analysis is as follows. All candidate muons inside the DMJ are selected by requiring $p_{T}(\mu)>1~\text{GeV}$ and $|\eta_{\mu}|<3.0$, parent particles ($\pi_{d}$’s) with $R_{xy}$ within one of the four regions defined in Table 6, and muon pairs with $\Delta\theta>0.02$ radians. Muons satisfying these criteria are identified as displaced muons. We employ the anti-$k_{t}$ algorithm Cacciari:2008gp as implemented in FASTJET Cacciari:2011ma to cluster these muons into jets, referred to as DMJs, using a jet radius parameter of $R=0.5$. In our analysis, we require at least six muons per jet, which renders the DMJ searches effectively background-free Zhang:2021orr .
Figure 10: The main kinematic distributions for the DMJ analysis at the CEPC with three benchmark points, $m_{\pi_{d}}=0.32$, $0.7$, and $1$ GeV, corresponding to $\kappa=5$ and $m_{X}=1$ TeV. Top-left: The dark pion transverse decay length, $R_{xy}$; Top-right: Muon transverse momentum, $p_{T}(\mu)$; Bottom-left: The number of muons in each muon jet; Bottom-right: The opening angle $\Delta\theta$ between two muons originating from the same dark pion.
We then select the leading $p_{T}$ DMJ in each event and present in Fig. 10 the main kinematic distributions of its constituent muons and their mother dark pions, using three BPs with $m_{\pi_{d}}=0.32~\text{GeV}$, $0.7~\text{GeV}$, and $1~\text{GeV}$, corresponding to $\kappa=5$ and $m_{X}=1~\text{TeV}$. The $R_{xy}$ (top-left) and $p_{T}(\mu)$ (top-right) distributions follow trends similar to those observed at low-energy $e^{+}e^{-}$ colliders, but at the CEPC they extend to significantly higher values due to its much larger collision energy. Since the CEPC detector is segmented into four non-contiguous radial regions as shown in Table 6, the $R_{xy}$ distribution reflects this structure; for three BPs considered here, most dark pions are concentrated in the innermost region, namely the vertex detector. The calculation of $R_{xy}$ also includes Gaussian smearing effects, with the transverse position resolution for displaced vertices given by CEPCStudyGroup:2018ghi :
[…]
where $a=5$, $b=$10\text{\,}\mathrm{G}\mathrm{e}\mathrm{V}$\$, $p_{T}$ is the average transverse momentum of the track, and $\theta$ is its polar angle. The distribution of the number of muons inside the leading $p_{T}$ DMJ (bottom-left) illustrates the multiplicity of displaced muons. Across all three BPs, DMJs exhibit high-multiplicity features, with lighter $\pi_{d}$’s showing a multiplicity advantage. Such high-multiplicity signatures are distinctive and very difficult for SM backgrounds to mimic Zhang:2021orr . The opening angle $\Delta\theta$ between two muons originating from the same dark pion (bottom-right) increases from BP1 to BP3. This wider separation for heavier $\pi_{d}$ results from the larger $p_{T}$ carried by the muons. However, due to the finite spatial resolution of the detector, muon pairs with excessively small $\Delta\theta$ cannot be resolved and are therefore excluded from the $\pi_{d}$ reconstruction.
To further identify the signal signature, all events containing at least two DMJs, each comprising more than six displaced muons, are selected. The signal efficiency $\epsilon_{S}$ is then evaluated from the above event selection criteria. The expected signal yield is given by $N_{S}=\mathcal{L}\times\sigma(e^{+}e^{-}\to q_{d}\bar{q}_{d})\times\epsilon_{S}$. Under the background-free hypothesis, the $2\sigma$ exclusion limit is set at $N_{S}=3$, from which we derive the exclusion contour in the $(m_{X},\kappa)$ plane, as shown in Fig. 9.
When determining the upper exclusion boundary, the same challenge as in the Belle II analysis is encountered, and thus conservatively set $N_{S}=10$ for the CEPC upper limit estimation, while the lower boundary remains determined by the condition $N_{S}=3$. However, for $m_{\pi_{d}}=0.32~\text{GeV}$, it remains hopeless to establish an upper boundary under this conservative assumption. We have estimated that more than $10^{9}$ Monte Carlo events are required for this situation beyond the scope of this work. Therefore, no upper exclusion boundary is presented for $m_{\pi_{d}}=0.32~\text{GeV}$.
VI Summary and discussion
We explored novel signatures of a leptophilic QCD-like composite asymmetric dark sector at $e^{+}e^{-}$ colliders by simulating dark parton showering, hadronization, and detector response to map the experimental reach. The study reveals a sharp transition in the dark pion proper decay length at the muon threshold (Fig. 2). Recast constraints from mono-photon searches and from slepton/stau reinterpretations limit the low-mass bi-charged scalar mediator region (Figs. 3 and 4), while combined projections for BaBar, Belle II, L-GAZELLE, and CEPC produce the characteristic band-shaped sensitivity in the $(m_{X},\kappa)$ plane as summarized in Fig. 11. After applying cosmological constraints on the $\pi_{d}$ lifetime, nearly all parameter space for $10~\text{MeV}\leq m_{\pi_{d}}<2m_{\mu}$ and $m_{X}\lesssim 10~\text{TeV}$ is excluded. Additionally, for $2m_{\mu}\leq m_{\pi_{d}}<2m_{\tau}$, a significant fraction of parameter space with $\kappa\gtrsim 1$ and $m_{X}\lesssim 10~\text{TeV}$ is also excluded.
Figure 11: Combined exclusion limits for different dark pion mass ranges. The first panel shows the exclusion region for $10~\text{MeV}<m_{\pi_{d}}<2m_{\mu}$, while the remaining three panels show benchmark points $m_{\pi_{d}}=0.32$, $0.7$ and $1~\text{GeV}$ from LHC, mono-photon searches at LEP and CEPC, displaced vertex searches (BaBar, Belle II, L-GAZELLE) and DMJ search at CEPC. All four panels include the parameter space where the proper lifetime of $\pi_{d}$ exceeds 1 sec, conservatively evaluated at $m_{\pi_{d}}=0.2~\mathrm{GeV}$ for the light mass range.
The primary experimental targets we identify are: (i) sparse displaced dimuon vertices at B-factories (kinematics in Figs. 57 and 8; trigger and event selection conditions are summarized in Tables 2 and 3), which Belle II can probe with realistic triggers and event selections; (ii) far detectors (L-GAZELLE) that extend sensitivity to substantially longer lifetimes; and (iii) displaced muon jet (DMJ) searches at high-energy $e^{+}e^{-}$ colliders (main kinematics shown in Fig. 10 and detector assumptions in Table 6). Although our projections adopt simplifying assumptions, lepton-flavor universality (LFU) of the mediator coupling, the benchmark relation $f_{\pi_{d}}=m_{\pi_{d}}$, and optimistic/background-free hypotheses in some regions, they nevertheless provide concrete, experimentally accessible targets and strongly motivate a coordinated program across B-factories, far detectors, and future $e^{+}e^{-}$ colliders.
Finally, we outline promising directions for future work. High-energy muon colliders could directly produce the bi-charged scalar mediator via Drell–Yan processes and thereby provide an independent probe of the parameter space. Advanced data-scouting or trigger-level analyses for low-$p_{T}$ displaced dimuon pairs at CMS and LHCb could significantly strengthen LHC constraints LHCb:2020ysn ; CMS:2023hwl ; Born:2023vll . When the dark pion is heavier than $2m_{\tau}$, dark showers become $\tau$-rich; this regime will require dedicated strategies (e.g., displaced and boosted $\tau$-tagging, and optimized reconstruction of hadronic $\tau$ decays) to retain sensitivity.
Acknowledgments
We thank Mengchao Zhang and Zeren Simon Wang for helpful discussions. CTL and CBX are supported by the National Natural Science Foundation of China (NNSFC) under grant No. 12335005, No. 12575118, and the Special funds for postdoctoral overseas recruitment, Ministry of Education of China. The authors gratefully acknowledge the valuable discussions and insights provided by the members of the China Collaboration of Precision Testing and New Physics.